# Noise-to-Notes: Diffusion-based Generation and Refinement for Automatic Drum Transcription

### v1

@@ 1 Introduction
Automatic drum transcription (ADT) is a subset of automatic music transcription, where the task is to derive symbolic musical representations from audio recordings. ADT primarily focuses on transcribing drum onset times, although the inclusion of velocity (dynamics) has also been shown to significantly enhance perceptual quality [1]. Despite the development of automatic methods to handle drum audio mixed with percussive or melodic instruments, consistent performance has only been achieved for simplified drum configurations [2]. Discriminating between different drum components is challenging because spectrograms derived from drum audio lack clear harmonic structure, with significant overlap in time and frequency components among instruments [2]. Moreover, the spectral characteristics of the same drum component may vary significantly depending on the sound sources and production methods used.
@@ 2 Related works
Leveraging rhythm game annotations, Zehren et al. [7] created the ADTOF dataset, containing 114 hours of real annotated drum onsets, and used this to train a CRNN to achieve state-of-the-art results across multiple ADT benchmarks. In addition to drum onsets, the E-GMD dataset [1] contains velocity annotations, and the associated Onset and Frames (OaF) Drums model, also based on CRNNs, achieved significantly better perceptual quality in human studies due to the inclusion of estimated velocities.
While CRNNs remain the most successful approach in ADT, transformers offer more flexible time and frequency modeling and have become the preferred approach for the closely related automatic piano transcription task [9, 10, 11]. Beyond discriminative models, there has also been recent interest in modeling piano transcription as a generative task using diffusion models [12, 13]. Diffusion models [14, 15] are a class of deep generative models that iteratively add noise to data and learn to reverse this process, enabling the generation of high-quality samples by progressively denoising from a stationary distribution. Diffusion models have been applied widely in music-related tasks, including music generation [16], editing [17] and restoration [18].
In the context of automatic piano transcription, DiffRoll [12] was the first method to reframe music transcription as a generative task, and used a U-Net-based diffusion model to predict piano onsets from mel-spectrogram features. More recently, DR3M [13] used a discrete denoising process with neigborhood attention to avoid modeling onsets as continuous variables, and trained an acoustic encoder to extract multi-scale mel-spectrogram features.
While modeling music transcription as a generative task has opened new possibilities for generation and refinement, neither DiffRoll nor DR3M reach the performance of discriminative models. Moreover, these methods only model onset, and the generative quality is unsatisfactory. Here, we introduce Noise-to-Notes (N2N), the first generative model for ADT. We propose a novel diffusion objective to facilitate joint onset-velocity prediction, demonstrate strong inpainting capability and improve robustness by integrating spectrogram and music foundation model (MFM) features, outperforming discriminative models and achieving state-of-the-art results across ADT benchmarks.
@@ 3 Method
Drum transcription is formulated as a frame-wise classification task, with the target transcription represented as $\textbf{x}\in\mathbb{R}^{N\times D\times 2}$, where $N$ is the number of frames and $D$ is the number of drum components. For each instrument at each frame, the ground truth annotations ($\text{{x}}_{0}$) consist of an onset $\in\{0,1\}$, indicating the presence of a drum hit, and associated velocities $\in[0,127]$ representing intensity. For training, we normalize both onset and velocity to [1,1][-1,1].
Here, we reframe ADT in a diffusion framework, where a denoising network $\mathcal{D}_{\theta}(\text{{x}}_{\sigma_{t}};\sigma_{t},\phi)$ is trained to recover a clean transcription from noisy transcriptions, conditioned on audio features ($\phi$). Concretely, the forward process $p(\text{{x}}_{\sigma_{t}};\sigma_{t})$ is used to generate noisy targets by adding i.i.d. Gaussian noise ($\sigma$) to a clean transcription from a predefined, monotonically increasing sequence of noise levels $\sigma_{0}=\sigma_{\text{max}}>\sigma_{1}>...>\sigma_{N}=\sigma_{\text{min}}$, such that the endpoint of the sequence corresponds to sampling from Gaussian noise at $\sigma_{\text{max}}$. During training, a randomly sampled noise level $t$ is used to generate noisy transcriptions $\text{{x}}_{t}\sim p(\text{{x}}_{t};\sigma_{t})$. The denoising network can then be trained by minimizing the mean-squared error (MSE) loss:
@@
However, modeling onsets and velocities as continuous values poses an optimization challenge and we observe that using this standard denoising objective is suboptimal (Sec 4.5). To avoid onset errors dominating the loss, we leverage the Pseudo-Huber loss [19], which has been used for training consistency models [20] and offers a smooth interpolation between MSE and MAE objectives using a fixed constant $c$. However, we observe that fixing $c$ hinders optimization (Sec 4.5), and instead we formulate an Annealed Pseudo-Huber loss ($\mathcal{L}_{\text{APH}}$):
@@
where $c(t)$ is scheduled during training to shift the objective from MSE loss at the beginning of training, to mean absolute error (MAE) loss near the end of training. Empirically, we use $c(t)=(1-\alpha_{t})c_{\text{max}}+\alpha_{t}c_{\text{min}}$, where $c_{\text{max}}=1$ and $c_{\text{min}}=1e^{-4}$.
Fig. 2: Noise-to-Notes architecture. N2N is an audio-conditioned transformer-based diffusion model. Drum audio features are extracted with a log mel-spectrogram and music foundation model. These features, combined with timestep (t) information, modulate the network through cross attention and FiLM layers.
Our proposed denoising network is shown in Fig. 2. We leverage the EDGE [21] architecture, a transformer-based decoder that incorporates timestep and music conditioning with feature-wise linear modulation (FiLM) [22]. For the audio condition, while mel-spectrogram features are commonly used in music transcription, intermediate features from MFMs may provide complementary information [23]. We leverage MERT [24] for the MFM, and extract both intermediate MFM and mel spectrogram features. Importantly, we apply two forms of dropout to both spectrogram and MFM features. To enable inpainting, we apply partial dropout, where a random, contiguous subsequence of the features are dropped. Moreover, we apply complete dropout, where the entire feature is dropped, to facilitate unconditional generation. To avoid confusion between the absence of noise and dropped-out features, we replace dropped-out regions with a learned null embedding. Moreover, we observed that applying a higher dropout rate to the spectrogram features is important to prevent overfitting. The features are then processed through separate transformer layers before being injected into the decoder to generate the denoised output. This output can be further refined by the reintroduction of Gaussian noise with a lower variance ($\text{{x}}_{t}\sim q(\text{{x}}_{t^{\prime}}|\text{{x}}_{t})$), and subsequent denoising at a lower timestep.
@@ 4.1 Implementation details
We trained our models using 4 NVIDIA A100 GPUs for $100$ epochs, with a learning rate of $3e^{-4}$ and batch size of $64$, taking $\approx 1$ day to complete. Besides using MERT as the audio encoder, we largely followed the same model settings as [21]. For log mel-spectrogram conditioning, to enable fair comparisons, we followed previous methods [1, 9], using an audio sample rate of 44.1kHz and hop length of $441$, resulting in 1010ms frames and 128 bins. For the MFM, we extracted intermediate (layer $10$) features from MERT [24] 330M model. We applied complete dropout with $p=0.15$ and $p=0.30$ for MERT and spectrogram features, respectively, and partial dropout with $p=0.5$ for both. We followed similar diffusion settings to EDM [25]. Specifically, we used a lognormal noise schedule for training: $\text{ln}(\sigma)\sim N(P_{\text{mean}},P_{\text{std}}^{2})$, and define noise scales for sampling:
$\left(\sigma_{\text{max}}^{\frac{1}{\rho}}+\frac{i}{N-1}\left(\sigma_{\text{min}}^{\frac{1}{\rho}}-\sigma_{\text{max}}^{\frac{1}{\rho}}\right)\right)^{\rho}.$ (3)
Moreover, we applied preconditioning to ensure that the magnitudes of the input and output remain consistent across different noise levels:
$\mathcal{D}_{\theta}(\text{x}_{\sigma_{t}},\sigma_{t},\phi)=c_{\text{out}}(\sigma_{t})F_{\theta}(c_{\text{in}}(\sigma_{t})\text{x}_{\sigma_{t}},c_{\text{noise}}(\sigma_{t}))+c_{\text{skip}}(\sigma_{t})\text{x}_{\sigma_{t}}.$ (4)
We used the same default hyperparameters as [25], where the definitions for $c_{\text{in}}$, $c_{\text{out}}$ and $c_{\text{noise}}$ can be found.
For evaluation, we used the music information retrieval library mir_eval [26] to compute note-wise onset and velocity transcription F1 scores. We followed established settings [2, 7, 1], assigning predicted note durations as 100100ms and used a tolerance of 5050ms.
4.2 Datasets
For training, we used the train partition of the E-GMD [1] dataset, containing 440 hours of human drum performances from 43 drum kits and the only publicly available dataset with associated onset and velocity MIDI annotations. For evaluation, we used the E-GMD test set, as well as the IDMT-SMT-Drums (IDMT) [27] and Medley DB Drums (MDB) [28] datasets. We used the same drum hit settings as [1], predicting 7 drum components (kick, snare, tom, hi-hat, crash cymbal, ride cymbal and bell). For MDB, we remapped drum outputs to a 5-hit setting [7], and for IDMT we used a 3-hit setting [1].
@@ 4.3 Evaluation
Unlike discriminative models, N2N uses a diffusion process that enables progressive refinement of transcriptions (Fig. 4). We observe the most dramatic performance improvements increasing one-step to few-step sampling. After around 10 sampling steps, we observe performance saturation. This suggests that the most significant performance improvements can be achieved with only a small increase in the inference costs, although it should be noted that there is an inference gap with discriminative models even at one-step (N2N: 0.163s, hFT-Transformer: 0.086s), due to the larger model size ($50M>5.5M$) and additional MFM feature extraction.
4.4 Generative Capability
Fig. 5: Inpainting example. Given a five-second clip where the last three seconds are masked, N2N uses available context to generate consistent transcriptions. KD=Kick. SN=Snare. HH=Hi-hat.
4.5 Ablation study
We performed an ablation study for our proposed methods (Table 2). Firstly, we observed that using the standard MSE loss function compromises velocity prediction performance. This is because onset errors dominated the overall loss, exacerbated by the squared error penalization. Using the original Pseudo-Huber loss improves velocity ($68.60>66.14$) but at the cost of onset ($80.11<85.55$). Using our proposed APH loss, there is a significant improvement in velocity ($76.17>66.14$), and even a moderate improvement to the onset ($86.77>85.55$), enabling effective learning of onset and velocity. We also investigated the effect of different audio features. While performance on E-GMD is similar whether spectrogram or MFM features are used, we observed dramatic improvements on the MDB ($82.16>71.15$) and IDMT ($90.36>80.89$) datasets using MFM features. To understand why conditioning on MFM features improves performance on out-of-domain data, we visualized spectrogram and MFM features from different drum datasets (Fig. 6). The MFM features between datasets are more distinct, supporting that these features contain high-level semantic information. This information can be useful to distinguish drum instruments from different domains that may have similar spectral characteristics. In fact, we observed that MFM features are complementary to spectrogram features, and obtained the best performance when both are used.
@@ 4.5 Ablation study
Fig. 6: t-SNE plots of spectrogram and MFM features. Spectrogram (left) and MFM (right) features were extracted from N2N input projection layer.
@@ 5 Conclusion
We reframed ADT as a generative task and proposed N2N, a diffusion model capable of transcribing drum onset and velocities from audio. We showed that using the standard diffusion objective is unable to optimize for both onset and velocity, and instead proposed an Annealed Pseudo-Huber loss that enables effective joint optimization. Moreover, we demonstrated the benefit of incorporating MFM features, in addition to commonly used spectrogram features, especially for improving transcription robustness on out-of-domain drum audio. Finally, we illustrated that reframing ADT into a generative task using diffusion models enables new inpainting and refinement capabilities. N2N sets a new state-of-the-art performance across multiple ADT benchmarks, and to our knowledge, establishes for the first time that generative models can surpass the performance of discriminative models for music transcription. In future work, we will investigate methods to bridge the inference gap to discriminative models, for example leveraging distillation to reduce model size and consistency modeling to improve sampling efficiency.

### v2

@@ 1 Introduction
Automatic drum transcription (ADT) is a subset of automatic music transcription, where the task is to derive symbolic musical representations from audio recordings. ADT primarily focuses on transcribing drum onset times, although the inclusion of velocity (dynamics) has also been shown to significantly enhance perceptual quality [1]. Despite the development of automatic methods to handle drum audio mixed with percussive or melodic instruments, consistent performance has only been achieved for simplified drum configurations [24]. Discriminating between different drum components is challenging because spectrograms derived from drum audio lack clear harmonic structure, with significant overlap in time and frequency components among instruments [24]. Moreover, the spectral characteristics of the same drum component may vary significantly depending on the sound sources and production methods used. Rather than directly using spectrogram features, music foundation models (MFMs) have been shown to provide effective features for various downstream music-related tasks [9] and this could be leveraged to disentangle the signals from different drum components.
While generative models have not been explored in ADT, recent research in other classification tasks suggests that they can outperform discriminative approaches [6]. The additional modeling of the data distribution could help mitigate overfitting, a significant issue in ADT. Here, we introduce the first generative model for ADT, Noise-to-Notes (N2N). We propose a novel diffusion objective to facilitate joint onset-velocity prediction, demonstrate strong inpainting capability and improve robustness by integrating spectrogram and MFM features, outperforming discriminative models and achieving state-of-the-art results across ADT benchmarks.
@@ 2 Related works
Leveraging rhythm game annotations, Zehren et al. [26] created the ADTOF dataset, containing 114 hours of real annotated drum onsets and trained a CRNN to achieve state-of-the-art results across multiple ADT benchmarks. In addition to drum onsets, the E-GMD dataset [1] contains velocity annotations, and the associated Onset and Frames (OaF) Drums model, also based on CRNNs, achieved significantly better perceptual quality in human studies due to the inclusion of estimated velocities. While CRNNs remain the most successful approach in ADT, transformers offer more flexible time and frequency modeling and have become the preferred approach for the closely related automatic piano transcription task [17, 25, 21].
Beyond discriminative models, there has also been recent interest in modeling piano transcription as a generative task using diffusion models [3, 8]. Diffusion models [15, 5] are a class of deep generative models that iteratively add noise to data and learn to reverse this process, enabling the generation of high-quality samples by progressively denoising from a stationary distribution. DiffRoll [3] was the first method to reframe music transcription as a generative task, and used a U-Net-based diffusion model to predict piano onsets from mel-spectrogram features. More recently, DR3M [8] used a discrete denoising process with neigborhood attention to avoid modeling onsets as continuous variables, and trained an acoustic encoder to extract multi-scale mel-spectrogram features.
While modeling music transcription as a generative task has opened new possibilities for generation and refinement, neither DiffRoll nor DR3M reach the performance of discriminative models. Moreover, these methods focus on modeling onset, and the generative quality remains limited.
@@ 3 Method
Drum transcription is formulated as a frame-wise classification task, with the target transcription represented as $\textbf{x}\in\mathbb{R}^{F\times D\times 2}$, where $F$ is the number of frames and $D$ is the number of drum components. For each instrument at each frame, the ground truth annotations ($\text{{x}}_{0}$) consist of an onset $\in\{0,1\}$, indicating the presence of a drum hit, and associated velocities $\in[0,127]$ representing intensity. For training, we rescale both onset and velocity to [1,1][-1,1].
Here, we reframe ADT in a diffusion framework, where a denoising network $\mathcal{D}_{\theta}(\text{{x}}_{\sigma_{t}};\sigma_{t},\phi)$ is trained to recover a clean transcription from noisy transcriptions, conditioned on audio and timestep features ($\phi$). Concretely, using the original DDPM formulation [5], the forward process $p(\text{{x}}_{\sigma_{t}};\sigma_{t})$ is used to generate noisy targets by adding i.i.d. Gaussian noise ($\sigma$) to a clean transcription from a predefined, monotonically increasing sequence of noise levels $\sigma_{0}=\sigma_{\text{max}}>\sigma_{1}>...>\sigma_{N}=\sigma_{\text{min}}$, such that the endpoint of the sequence corresponds to sampling from Gaussian noise at $\sigma_{\text{max}}$. During training, a randomly sampled noise level $t$ is used to generate noisy transcriptions $\text{{x}}_{t}\sim p(\text{{x}}_{t};\sigma_{t})$. The denoising network can then be trained by minimizing the mean-squared error (MSE) loss:
@@
However, modeling onsets and velocities as continuous values poses an optimization challenge and we observe that using this standard denoising objective is suboptimal (Sec 4.3). To avoid onset errors dominating the loss, we leverage the Pseudo-Huber loss [2], which has been used for training consistency models [14] and offers a smooth interpolation between MSE and mean absolute error (MAE) objectives using a fixed constant $c$. However, we observe that fixing $c$ hinders optimization (Sec 4.3), and instead we formulate an Annealed Pseudo-Huber loss ($\mathcal{L}_{\text{APH}}$):
@@
where $c(t)$ is scheduled during training to shift the objective from MSE loss at the beginning, to MAE loss near the end of training. Empirically, we use $c(t)=(1-\alpha_{t})c_{\text{max}}+\alpha_{t}c_{\text{min}}$, where $c_{\text{max}}=1$, $c_{\text{min}}=10^{-4}$, and $\alpha_{t}$ is linearly annealed from 0 to 1.
Fig. 2: Noise-to-Notes architecture. N2N is an audio-conditioned transformer-based diffusion model. Drum audio features are extracted with a log mel-spectrogram and music foundation model. These features, combined with timestep ($\sigma_{t}$) information, modulate the decoder through cross attention and FiLM layers.
In practice, we used a more recent diffusion framework [7] that includes several modifications to improve generation efficiency and quality. Specifically, a lognormal noise schedule is used for training: $\ln\sigma\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$, with defined noise scales for sampling:
$\left(\sigma_{\max}^{1/\rho}+\frac{i}{N-1}\left(\sigma_{\min}^{1/\rho}-\sigma_{\max}^{1/\rho}\right)\right)^{\rho}.$ (3)
For our experiments, we used the same default hyperparameters as [7].
Our proposed denoising network is shown in Fig. 2. We leverage the EDGE [18] architecture, a transformer-based decoder that incorporates timestep and music conditioning with feature-wise linear modulation (FiLM) [12]. For the audio condition, while mel-spectrogram features are commonly used in music transcription, intermediate features from MFMs may provide complementary information [10]. We leverage MERT [9] for the MFM, and extract both mel spectrogram and intermediate MFM features for the encoder. Importantly, we apply two forms of dropout to both spectrogram and MFM features. To enable inpainting, we apply partial dropout, where a random, contiguous subsequence of the features are dropped. Moreover, we apply complete dropout, where the entire feature is dropped, to facilitate unconditional generation. To avoid confusion between the absence of noise and dropped-out features, we replace dropped-out regions with a learned null embedding. Moreover, we observed that applying a higher dropout rate to the spectrogram features is important to prevent overfitting. The features are then processed through separate transformer layers before being injected into the decoder to generate the denoised output. This output can be further refined by the reintroduction of Gaussian noise with a lower variance ($\text{{x}}_{t}\sim q(\text{{x}}_{t^{\prime}}|\text{{x}}_{t})$), and subsequent denoising at a lower timestep.
@@ 4.1 Implementation details
Experimental settings. We trained our models using 4 NVIDIA A100 GPUs for $100$ epochs, with a learning rate of $3\times 10^{-4}$ and batch size of $64$, taking $\approx 1$ day to complete. Besides using MERT as the audio encoder, we largely followed the same model settings as [18]. For log mel-spectrogram conditioning, to enable fair comparisons, we followed previous methods [1, 17], using an audio sample rate of 44.1 kHz and hop length of $441$, resulting in $10$ ms frames and 128 bins. For the MFM, we extracted intermediate (layer $10$) features from MERT 330M model. We applied complete dropout with $p=0.15$ and $p=0.30$ for MERT and spectrogram features, respectively, and partial dropout with $p=0.5$ for both.
Datasets. For training, we used the train partition of the Expanded Groove MIDI Dataset (E-GMD) [1], containing 440 hours of human drum performances from 43 drum kits and the only publicly available dataset with associated onset and velocity MIDI annotations. For evaluation, we used the E-GMD test set, as well as the IDMT-SMT-Drums (IDMT) [4] and Medley DB Drums (MDB) [16] datasets. We used the same drum hit settings as [1], predicting 7 drum components (kick (KD), snare (SN), tom (TT), hi-hat (HH), crash cymbal (CY), ride cymbal (RD) and bell (BE)). For MDB, we remapped drum outputs to a 5-hit setting [26], and for IDMT we used a 3-hit setting [1].
Evaluation metrics. For evaluation, we used the music information retrieval library mir_eval [13] to compute note-wise onset and velocity transcription F1 scores. We followed established settings [24, 26, 1], assigning predicted note durations as 100100ms and used a tolerance of 5050ms.
@@ 4.2 Evaluation
Unlike discriminative models, N2N uses a diffusion process that enables progressive refinement of transcriptions (Fig. 4). We observe the most dramatic performance improvements increasing one-step to few-step sampling. After around 10 sampling steps, we observe performance saturation. This suggests that the most significant performance improvements can be achieved with only a small increase in the inference costs, although it should be noted that there is an inference gap with discriminative models even at one-step (N2N: 0.163 s, hFT-Transformer: 0.086 s), due to the larger model size (5050M>>5.55.5M) and additional MFM feature extraction.
@@
Fig. 5: t-SNE plots of spectrogram (left) and MFM (right) features. Features were extracted from N2N input projection layer.
4.3 Ablation study
We performed an ablation study for our proposed methods (Table 2). We observed that the standard MSE loss function compromises velocity prediction performance, due to onset errors dominating the overall loss, exacerbated by the squared error penalization. Using the Pseudo-Huber loss improved velocity ($68.60>66.14$) but at the cost of onset ($80.11<85.55$). Using our proposed APH loss, there is a significant improvement in velocity ($76.17>66.14$), and even a moderate improvement to the onset ($86.77>85.55$). We also investigated the effect of different audio features. While performance on in-domain E-GMD is similar, we observed dramatic improvements on the MDB ($82.16>71.15$) and IDMT ($90.36>80.89$) datasets using MFM features. To understand why conditioning on MFM features improves performance on out-of-domain data, we visualized spectrogram and MFM features from different drum datasets using a two-dimensional t-SNE [11] (Fig. 5). The MFM features between datasets are more distinct, supporting that these features contain high-level semantic information. This information can be useful to distinguish drum instruments from different domains that may have similar spectral characteristics. In fact, we observed that MFM features are complementary to spectrogram features, and obtained the best performance when both are used.
Fig. 6: Inpainting example. Given a five-second clip where the last three seconds are masked (replaced with null embedding), N2N uses available context to generate consistent transcriptions.
4.4 Generative Capability
@@ 5 Conclusion
We reframed automatic drum transcription (ADT) as a generative task and proposed Noise-to-Notes (N2N), a diffusion model capable of transcribing drum onset and velocities from audio. We showed that using the standard diffusion objective is unable to optimize for both onset and velocity, and instead proposed an Annealed Pseudo-Huber loss that enables effective joint optimization. Moreover, we demonstrated the benefit of incorporating MFM features, in addition to commonly used spectrogram features, especially for improving transcription robustness on out-of-domain drum audio. Finally, we illustrated that reframing ADT into a generative task using diffusion models enables new inpainting and refinement capabilities. N2N sets a new state-of-the-art performance across multiple ADT benchmarks, and to our knowledge, establishes for the first time that generative models can surpass the performance of discriminative models for automatic music transcription. In future work, we will investigate methods to bridge the inference gap to discriminative models, for example leveraging distillation to reduce model size and consistency modeling to improve sampling efficiency. Additionally, we aim to investigate extending this approach to address multi-instrument transcription.